### v1

@@ Abstract
Involving people in energy systems planning can increase the legitimacy and socio-political feasibility of energy transitions. Participatory research in energy modelling offers the opportunity to engage with stakeholders in a comprehensive way, but is limited by how results can be generated and presented without imposing assumptions and discrete scenarios on the participants. To this end, we present a methodology and a framework, based on near-optimal modelling results, that can incorporate stakeholders in a holistic and engaging way. We confront stakeholders with a continuum of modelling-based energy system designs via an interactive interface allowing them to choose essentially any combination of components that meet the system requirements. Together with information on the implications of different technologies, it is possible to assess how participants prioritise different aspects in energy systems planning while also facilitating learning in an engaging and stimulating way. We showcase the methodology for the remote Arctic settlement of Longyearbyen and illustrate how participants deviate consistently from the cost optimum. At the same time, they manage to balance different priorities such as emissions, costs, and system vulnerability leading to a better understanding of the complexity and intertwined nature of decisions.
@@ 1 Introduction
Our transition towards more sustainable energy systems needs to accelerate, but its substantial challenges should not be underestimated. Such transformations require considerable technological change, development and innovation, but are also increasingly recognised to rely on changes in human practices [1] as well as societal and citizen engagement [2, 3]. Despite the fact that social acceptance and political feasibility are considerable barriers to energy system transitions [4], much research analysing future energy systems devotes surprisingly little focus to it [5, 6, 7, 8]. Energy transitions involving people in decision making creates legitimacy of the processes, enables change and can help to ensure socially just outcomes.
Participatory research in energy modelling can fill this gap, but lacks methodologies that are holistic and engaging at the same time. We argue in this article that a lot can be gained for the field and the public discourse if these methodologies lay common ground in a realistic and target-oriented fashion. To this end we introduce an interactive framework between policymakers, modellers, and stakeholders in which near-optimality helps to assess socio-political priorities in the energy transition and demonstrate that it facilitates learning at the same time.
Despite a generally favourable public perception of renewable electricity technologies [9, 10] and transitioning to low-carbon energy systems altogether in Europe, there are diverse perceptions and opinions on the favoured end-state and pathways to get there [11]. Transitioning to low-carbon energy systems will naturally come with both social and environmental consequences. It may lead to an unequal distribution of air pollution [12], job creation and losses [13] and visual impact [14]. Additionally, terrestrial and bird species may suffer from habitat loss, conversion or fragmentation following the construction of hydropower infrastructure, power lines or wind power plants [15, 16]. Understanding the impact of such factors on public acceptance and support is critical for energy systems planning.
Although surveys are commonly used to query both laypeople and experts, they may not be best suited to evaluate different energy system components: the knowledge gaps about the technologies in question can lead to misconceptions of risks and impacts [17, 18]. This can be alleviated by approaches aiming to elicit informed public preferences such as factsheets [19], group discussions [17], and interactive tools [19, 20]. Such tools and information material have been shown to generate more stable preferences [18], create more knowledge and are additionally appreciated by participants [19, 21, 17]. A consistent methodological challenge in all the above techniques is that of representing the underlying energy system holistically, capturing the complex interplay of variable renewables, transmission and storage [22].
Energy system models are important tools used to study energy systems and transitions, and their techno-economic nature gives insights of different transition pathways, sensitivity to developments and point out which challenges exist. Increasingly, energy modellers engage in participatory modelling: they involve different stakeholders already in the modelling process through scenario generation, surveys and interviews, or multi-criteria decision making [23]. However, only about one in six of previously reviewed participatory modelling studies included non-academic experts [23], indicating that particular stakeholder groups may be overlooked.
Participatory approaches to energy systems modelling are seen as important ways to generate bottom-up knowledge of assumptions and narratives [24, 23], assess which designs would be socially accepted [25] and to facilitate mutual learning for both researchers and involved actors [23]. At the same time, key challenges include dealing with the complexity of the models as well as transparency of assumptions and how the models work [23]. Presenting stakeholders with results based on a few scenarios and criteria may constrain the diversity of options available, potentially based on normative preconceptions and assumptions by researchers. Instead, modelling frameworks in which stakeholders themselves are allowed to define metrics and explore alternative system configurations are preferred [24, 26].
One method used in the literature which allows for this is to explore the near-optimal feasible space of a model, using modelling to generate alternatives (MGA) [27]. Beyond only exploring cost-optimal modelling, MGA allows modellers to deviate from the cost optimum and systematically explore alternatives which may be slightly more costly, but that score better on other non-modelled objectives, such as political feasibility, energy security concerns or distributional outcomes [28, 29]. Whereas MGA has been used to explore energy systems that are robust to weather uncertainty [30], effort-sharing schemes and inequity [31] as well as flexibility and trade-offs from different system configurations [32, 33, 34], its applicability for stakeholder involvement and participatory modelling is yet to be explored in practice. Near-optimal feasible spaces present an opportunity to better draw on the holistic perspectives of energy system models when engaging with stakeholders to understand the underlying reasons for preferences as well as for their use as information material.
@@
With this work, we investigate how participatory modelling informed by near-optimal solutions can inform 1) stakeholders about transition possibilities and 2) policymakers about stakeholder’s priorities. We demonstrate the methodology for the remote Arctic settlement of Longyearbyen, Svalbard, due to its remote location and inability to externalise the impact of energy system infrastructure. We derive a large sample of energy system configurations, based on the results from an energy systems optimisation model (PyPSA-LYB), that are all feasible and which represent the diversity of options for the local energy system. These energy system configurations are then presented to local inhabitants of Longyearbyen through an interactive web interface, with which we study the technological preferences, priorities and rationale of the participants. The uniqueness of this interactive interface is that it combines “real” energy system designs that cover the interactions and system aspects of an energy system model with explicit impacts. Our implementation, which dynamically updates selectable design parameters in order to restrict the user to only feasible system designs, represents a novel and accessible way to explore and interact with the vast space of options generated by near-optimal methods. Despite being used in a specific geographical context, the conceptual framework and tool can be implemented in other settings.
@@ 2 A new modelling-based interactive planning tool
Figure 1: Overview over Longyearbyen and its current energy system. On the left, the map shows the geographical location of Longyearbyen on Svalbard as well as its distance from the mainland ($\sim$950 km) making an electricity link prohibitively expensive. On the top right, the per capita greenhouse gas emissions for Longyearbyen (and Norway as comparison) are plotted: for the coal-powered CHP power plant (until 2022), and the current diesel-powered CHP power plant [35, 36]. Norwegian emissions are computed based on 2021 data on electricity consumption (https://www.ssb.no/statbank/table/06913/) and its carbon density (https://www.nve.no/energi/energisystem/kraftproduksjon/hvor-kommer-stroemmen-fra/) as well as biomass heating emissions from Arvesen et al. [37]. On the bottom right, the monthly average capacity factors for onshore wind and solar PV are plotted (for the input weather year). Note that these are subject to interannual variability and therefore just one realisation.
Longyearbyen (see figure 1 for an overview of the system) is well-suited for energy systems research for a number of reasons. For one, the energy system of Longyearbyen is an isolated off-grid system, located at 78°N on the Arctic archipelago of Svalbard, and connecting with the north of Norway ($\sim$950km away) is considered expensive and unrealistic [38]. The external influences on the system are therefore limited, which simplifies modelling, and at the same time make impacts and trade-offs of different energy technologies directly apparent. Longyearbyen is already transitioning away from its coal-powered history and is of extraordinary geopolitical interest [39] — which keep energy on the agenda of the local population, and is frequently discussed in local and national media.
@@
We presented an interactive interface that allows the local community to participate in energy planning. Although designed to engage the inhabitants of Longyearbyen, the methods and open-source code we develop also lend itself to other contexts. Figure 2 illustrates both the generalised (left) and particular (right) workflow used. We engaged with the local population in public spaces between March 11 and 15, 2024; each participant interacted with the tool, based on results from the energy system optimisation model PyPSA-LYB (Methods). The scenarios underlying the model have previously been used for planning the energy transition in Longyearbyen [40] and serve as a baseline for our modelling development. With the transition plan [40] defining the limits and broad scope of modelling assumptions, we explore near-optimal solutions (also known as modelling-to-generate-alternatives (MGA)) to paint a realistic picture of how alternative energy system designs for Longyearbyen could look like. In particular, we follow Grochowicz et al. [30] to generate the near-optimal feasible space and sample 56,050 different energy system configurations (see Methods for extensive details).
@@
In the interface, participants explored a continuum of energy system designs, by allocating investment in different system components. Within the feasible bounds (pre-computed in PyPSA-LYB) and a generous cost increase (of up to 125% on top of the optimal costs), participants were free to set investment levels in onshore wind power, solar power, imported green fuels, heat storage and hydrogen infrastructure (see Methods).
@@
Figure 3: Overview of the main page of the user interface. The sliders and decision-variables which users may change is shown in a), whereas b) shows the resulting effects of the chosen system configuration. The interface also include a feature (button) to minimise a specific metric, which participants could use as a starting point. Additionally, the cost-optimal configuration is shown with a vertical line for each slider. However, it was not the starting point when opening the interface.
@@
In practice, the level of investments is controlled through sliders, as seen in a) of figure 3 and restricted by an updated geometry of the near-optimal space relative to the slider location. Apart from the investment levels, the participants were presented with key metrics of their chosen system configuration. These metrics are computed heuristically as needed by the interface, interpolating between pre-computed values for the 56,050 sampled configurations (Methods). This feature, illustrated in b) of Figure 3, adds relevant information and context to the decision process, and keeps track of our research question aimed at (balancing of) priorities of local stakeholders. By complementing these quantitative values to possible qualitative pre-conceptions, the participants saw the impacts of their choices directly and could identify an energy system design best representing their preferences. After submitting their final energy system designs, participants were also asked to choose up to three of the metrics (some chose more than three) that they prioritised for the energy system design — i.e. their stated priorities — how much additional cost they were willing to accept to improve the representation of their priorities and provide written feedback to the process. Complete details on the interface, as well as the rationale for choosing the particular dimensions and metrics are described in Methods.
@@ 3 Cost-optimality paradigms cannot meet diverse expectations and technological perspectives
The outputs from the interface showcase the variety of proposed system designs and give insights about the acceptance of technologies, the importance of specific metrics as well as a partial rebuttal of cost-optimal designs, the gold standard of energy system modelling. Through our exemplary work comprising 5% of the population (126 participants, see Methods for details on the demographics of the participants), we can also infer specific trends and patterns for system planning in Longyearbyen.
@@
Figure 4: Combined strip- and box plot of submitted results across the five dimensions. Stars represent the cost-optimal investment combination and one million NOK is approximately equivalent to 86,000 EUR or 90,500 USD.
@@
None of the technologies have consistent results, and the interquartile ranges (IQRs) are between 10.9 and 36.8 percent of the decision interval. This is somewhat expected, as the design of the interface as well as the high slack level (additional 125% of costs were allowed for solutions to remain near-optimal) provide the participants with great flexibility and a high degree of freedom to select their preferred combination of technologies. In addition to the 56,050 discrete energy system designs precomputed, the participants are also able to pick points anywhere within the feasible space, thanks to post-processing and interpolation happening in the interface (see Methods for details), further increasing the diversity of options.
@@
Some of the results, the median for wind power investment and import of green fuels for instance, are relatively close to the cost-optimal investment. Solar power and hydrogen storage are out-competed in the cost optimum, but we see that most of the participants have opted for considerable investment into these technologies. The min-max normalised median absolute deviation (MAD) quantify the distance between the median and the cost-optimum (as illustrated in figure 4) for the five dimensions, and are 0.06, 0.34, 0.11, 0.17 and 0.22, respectively. It is clear that the participants deviate from the cost-optimal solution, and that they are willing to accept additional costs, if it improves other, non-monetised objectives.
@@
The resulting energy system designs of the participants show very high slack values (mean 81% and median 91%), which raises questions about how well the additional cost is understood by the participants and who would bear that cost. This is striking given that almost half of the participants prioritised electricity prices when designing their system (more on that later). On the contrary, when the participants were asked how much additional cost they were willing to accept, the results were much lower (mean and median values of 26% and 15% respectively) in the post-study questionnaire. This stark divide between revealed and stated preferences hints at the challenge to make results more tangible, but also shows that participants expect high performance on many priorities including affordability simultaneously.
@@
4 Successful energy transitions rely on balancing trade-offs of local priorities
@@
While the aggregated results hint at the popularity of less cost-effective and less invasive technologies like solar power, heat and hydrogen storage, we can assess the participants’ choices based on their stated priorities. According to the post-study questionnaire, the most commonly prioritised metrics are emissions (69%), followed by vulnerability (49.2%) and electricity price (47%). The remaining four metrics were prioritised by considerably fewer people (visual impact — 25.4%, system costs — 24.6%, heat price — 21.4% and land use — 15.9%).
Figure 5: Output values depending on whether a metric was prioritised.
@@
With the flexibility and diversity of options available, participants are forced to balance different priorities and trade-offs. Most participants (57%) stated that they prioritised three or more metrics when making their decision, introducing complexity into their decision-making. In fact, many participants openly acknowledged that their attempts of multi-objective planning is what they expected from policymakers too. Figure 5 evaluates whether participants were successful in achieving their stated priorities when they used the tool: we compare the resulting metrics and the difference in outcome for each of the metrics, categorised by whether the participant stated to prioritise this metric when designing their system. For each of the metrics, except heat price, the group of participants who prioritised a metric, are, on average, able to identify system designs improving that metric, compared to those who did not. Some metrics, particularly costs, have a larger differences than other.
@@
Figure 6: Obtained values for vulnerability and emissions from participants’ choices. Each point in a) represent a participant, and their system’s vulnerability and emissions. The dashed outer line represent the complete near-optimal feasible space and the colour of a point indicate the resulting slack. The convex hulls for seven slack levels in b) show how the decision-space change with increasing slack.
@@
As already established, participants tried to balance multiple priorities, further complicating the evaluation of how well people manage to achieve their priorities (also given that their individual weighing of the different metrics is unknown to us). In a) of figure 6, we illustrate how the participants have balanced the two most prioritised metrics, vulnerability and emissions, and the resulting slack value through the shade of the point. Most of the participants have chosen systems with very low emissions, whereas there is more variability for vulnerability. Furthermore, b) of figure 6 show how the decision space of the participants grows with increasing slack values. The figure shows that it is possible to balance these two priorities, even at low slack levels, but further reducing the vulnerability come at the expense of a more expensive system.
@@
Figure 7: Correlation matrix between revealed slider choice and stated priority of participants
@@
The correlation matrix between stated priorities and revealed slider choices is shown in figure 7, which illustrates patterns between the choices of the participants and their stated priorities. Participant who prioritised additional costs generally reduced investment (and over-sizing of the system), whereas prioritising emissions correlate with high investment in solar power and vulnerability with additional investment into heat storage as well as wind and solar power. Prioritising visual impact naturally led to reduced investment in wind power (the culprit for visual impact), and instead showed a tendency for elevated investments in solar power and green imports. Participants who stated a high willingness to pay also leaned into heat storage investment.
@@
The participants were restricted to solutions that are feasible system configurations, and they were therefore unable to simply minimise all of their priorities, and instead forced to make trade-offs and balance their priorities. This challenged some of the participants to think about the complexity (R01, R02, R03 in table 1), but also contributed to a feeling of being more informed.
Furthermore, there were also participants who made the case for technologies and solutions that were not available in the interface, such as returning to a coal-powered energy supply (R04, R05), demand reducing measures (R06) and nuclear power (R07).

### v2

@@ Abstract
Including local populations in energy systems planning can increase the legitimacy and socio-political feasibility of energy transitions. Participatory modelling enables comprehensive involvement of stakeholders and can facilitate learning; however, the choice of options and scenarios in the model can be opaque to participants. Here we present a framework in which stakeholders engage with near-optimal modelling results in a holistic way, and showcase our methodology in a case study conducted in the remote Arctic settlement of Longyearbyen. Participants choose across a continuum of modelling-based energy system designs in an interactive interface through which they select essentially any feasible combination of system components, and immediately see the implications of their choice. We observe that stakeholders in Longyearbyen consistently deviate from the cost optimum and instead balance different aspects like emissions, costs, and system vulnerability. When confronted with difficult trade-offs through our tool, participants feel better informed and successfully navigate through complex and intertwined decision-making.
@@ 1 Introduction
Our transitions towards more sustainable energy systems need to accelerate, but face substantial challenges. Such transformations require considerable technological change, development, and innovation, but are also increasingly recognised to rely on changes in human practices [1] as well as societal and citizen engagement [2, 3]. Despite barriers of social acceptance and political feasibility [4], much research analysing future energy systems devotes surprisingly little focus to it [5, 6, 7, 8]. Energy transitions involving people in decision-making create legitimacy of the processes, enables change and can help ensure socially just outcomes [9].
Participatory research in energy modelling aims to fill this gap, but currently lacks holistic and simultaneously engaging methodologies. In this article, we introduce a multi-step framework through which policymakers, stakeholders, and modellers can learn from each other and interact to find more socially acceptable energy transitions; we demonstrate this framework with a case study conducted in the Arctic community of Longyearbyen. Our methodology is rooted in a vast number of near-optimal realistic system designs presented in an interactive interface, allowing study participants to consider trade-offs and communicate priorities for system planning. We introduce a modelling approach, to our knowledge the first of its kind, in which participants can interact with and decide from a vast space of feasible energy systems to design a system of their liking.
Despite generally favourable public perceptions of renewable electricity technologies [10, 11] and transitions to low-carbon energy systems altogether in Europe, opinions vary on the favoured end-state and pathways to get there [12]. The overhaul of energy infrastructure to low-carbon systems comes with environmental [13, 14] as well as social consequences [15, 16, 17]. Although it is crucial to consider such (in-)direct impacts for public acceptance, e.g. querying laypeople and experts, the knowledge gaps about technologies in question can lead to misconceptions of risks and impacts [18, 19] when surveys are used to evaluate different energy system components. Other informative approaches such as factsheets [20], group discussions [18], and interactive tools [20, 21] generate more stable preferences [19], create more knowledge and are additionally appreciated by participants [20, 22, 18]. Yet, in all of the above it is challenging to represent the underlying system holistically, capturing the complex interplay of variable renewables, transmission and storage [23]. This motivates the usage of energy system models whose techno-economic nature gives insights about different transition pathways and which challenges exist. Increasingly, energy modellers engage in participatory modelling: they involve stakeholders already at the modelling stage through scenario generation, surveys and interviews, or multi-criteria decision-making [24]. Particular stakeholder groups may still be overlooked, as only one in six reviewed participatory modelling studies included non-academic stakeholders [24]. Nonetheless, participatory modelling can generate bottom-up knowledge of assumptions and narratives [25, 24], help assess which designs would be socially accepted [26] and facilitate mutual learning for both researchers and involved actors [24]. With increasingly complex models, the success of such approaches hinges on transparency of assumptions and modelling tools [24], but presenting stakeholders with results based on a few pre-selected scenarios and criteria may constrain the diversity of options available, potentially based on normative preconceptions and assumptions by researchers. Instead, modelling frameworks in which stakeholders themselves are allowed to define metrics and explore alternative system configurations are preferred [25, 27].
A systematic exploration of alternative system designs is possible by describing the near-optimal feasible space of a model, also known as modelling-to-generate alternatives (MGA) [28]. Looking beyond only cost-optimal modelling, MGA finds feasible solutions that are only slightly more expensive, but score better on other objectives: it is possible to directly search for more diverse alternatives [29, 30], classify trade-offs [31], find designs that are robust to weather uncertainty [32], or perform better on non-modelled objectives such as inequity [33], political feasibility, energy security concerns or distributional outcomes [34, 35]. Despite the myriad applications of MGA, and the potential for communicating trade-offs and informing policymakers, it has not been applied to involve stakeholders in participatory modelling efforts. Here we bridge this gap by creating a framework through which near-optimal spaces describe the diversity of possible solutions which are accessed and communicated through an interactive interface.
@@
In this work, we investigate how participatory modelling by near-optimal solutions can inform 1) stakeholders about transition possibilities and 2) policymakers about stakeholder’s priorities. We demonstrate the methodology with a holistic view on the move away from coal in the remote Arctic settlement of Longyearbyen, Svalbard. Its remote location and inability to externalise the impact of energy system infrastructure makes it overseeable and particularly suitable for stakeholder engagement. We pre-compute 56,050 feasible energy system configurations with an energy system optimisation model (PyPSA-LYB) leading to diverse outcomes for the community which are accompanied by metrics measuring different qualities of each system design. These configurations are presented to inhabitants of Longyearbyen through a novel interactive interface which allows us to assess the technological preferences, priorities and rationale of the participants while ensuring that system requirements are met at all times. Our implementation, which dynamically updates selectable design parameters in order to restrict the user to only feasible system designs, represents a novel and accessible way to explore and interact with the vast space of options generated by near-optimal methods. Despite being used in a specific geographical context, the conceptual framework and tool can be implemented in other settings.
@@ 2 A new modelling-based interactive planning tool
Figure 1: Overview over Longyearbyen and its current energy system. On the left, the map shows the geographical location of Longyearbyen on the Arctic archipelago of Svalbard as well as its distance from the north of Norway ($\sim$950 km) making an electricity link prohibitively expensive and unrealistic. On the top right, the per capita greenhouse gas emissions for Longyearbyen (and Norway as comparison) are plotted: for the coal-powered CHP power plant (until 2022), and the current diesel-powered CHP power plant [36, 37]. Norwegian emissions are computed based on 2021 data on electricity consumption (https://www.ssb.no/statbank/table/06913/) and its carbon density (https://www.nve.no/energi/energisystem/kraftproduksjon/hvor-kommer-stroemmen-fra/) as well as biomass heating emissions from Arvesen et al. [38]. On the bottom right, the monthly average capacity factors for onshore wind and solar PV are plotted (for the input weather year). Note that these are subject to interannual variability and therefore just one realisation.
Longyearbyen (overview in figure 1) is well-suited for energy systems research for a number of reasons. For one, the energy system of Longyearbyen is an isolated off-grid system, located at 78°N on Svalbard, and a mainland connection ($\sim$950km away) is prohibitively expensive [39]. The external influences on the system are therefore limited, which simplifies modelling, and at the same time makes impacts and trade-offs of different energy technologies directly apparent. Longyearbyen is already transitioning away from its coal-powered history and is of extraordinary geopolitical interest [40] — which keeps energy on the agenda of the local population, and is frequently discussed in local and national media.
@@
To engage the local community in energy planning, we developed an interactive interface, which functions as a tool for users to interact with results from an energy system optimisation model, PyPSA-LYB (Methods). The scenarios underlying the model have previously been used for planning the energy transition in Longyearbyen [41] and serve as a baseline for our modelling development. With the transition plan [41] defining the limits and broad scope of modelling assumptions, we explore near-optimal solutions to paint a realistic picture of how alternative energy system designs for Longyearbyen could look like. In particular, we follow Grochowicz et al. [32] to generate the near-optimal feasible space and sample 56,050 different energy system configurations (see Methods).
@@
We engaged with the local population in public spaces between March 11 and 15, 2024; each participant interacting with the interface individually. In the interface, participants explored a continuum of energy system designs, by allocating investment in different system components. Within the feasible bounds (pre-computed in PyPSA-LYB) and a generous cost increase (of up to 125% on top of the optimal costs), participants freely set investment levels for onshore wind power, solar power, imported green fuels, heat storage and hydrogen infrastructure (see Methods). Figure 2 illustrates both the generalised (left) and particular (right) workflow used.
@@
Figure 3: Overview of the main page of the user interface. The sliders and decision-variables which users may change are shown in a), whereas b) shows the resulting effects of the chosen system configuration. The interface also includes a feature (buttons) which allows choosing one of the metrics to be minimised, which participants could use as a starting point. Additionally, the cost-optimal configuration is shown with a vertical line for each slider. However, it is not the starting point when opening the interface, which was chosen at random.
@@
In practice, the level of investments is controlled through sliders, as seen in figure 3a) and restricted by an updated geometry of the near-optimal space relative to the slider location. Apart from the investment levels, the participants were presented with key metrics of their chosen system configuration. These metrics are computed heuristically as needed by the interface, interpolating between pre-computed values for the 56,050 sampled configurations (Methods). This feature, illustrated in figure 3b), adds relevant information and context to the decision process, and keeps track of our research question aimed at (balancing) priorities of local stakeholders. By complementing these quantitative values with possible qualitative preconceptions, the participants saw the impacts of their choices directly and could identify an energy system design best representing their preferences. After submitting their final energy system design, participants were also asked to choose up to three metrics (some chose more than three) that they prioritised for the energy system design — i.e. their stated priorities —, how much additional cost they were willing to accept to improve the representation of their priorities and provide written feedback to the process. Complete details on the interface, as well as the rationale for choosing the particular dimensions and metrics are described in Methods.
@@ 3 Cost-optimality paradigms cannot meet diverse expectations and technological perspectives
The interface outputs showcase the variety of proposed system designs and give insights about the acceptance of technologies, the importance of specific metrics as well as a partial rebuttal of cost-optimal designs, the gold standard of energy system modelling. Through our exemplary work comprising 5% of the population (126 participants, demographics in Methods), we can also infer specific trends and patterns for system planning in Longyearbyen.
@@
Figure 4: Combined strip- and boxplot of submitted results across the five dimensions. Stars represent the cost-optimal investment combination and one million NOK is approximately equivalent to 86,000 EUR or 90,500 USD in 2025. The min-max normalised median absolute deviations (MAD) quantify the distance between the median and the cost-optimum for the five dimensions, and are 0.06, 0.34, 0.11, 0.17 and 0.22, respectively.
@@
None of the technologies have consistent results with interquartile ranges (IQRs) of 10.9–36.8 percent of the decision interval. This is somewhat expected, as the design of the interface as well as the high slack level (125% more expensive systems are “near-optimal”) provide the participants with a high degree of freedom to select their preferred combination of technologies. In addition to the pre-computed 56,050 discrete energy system designs, the participants are also able to pick points anywhere within the feasible space, as the post-processing and interpolation within the interface (Methods) increases the diversity of options.
@@
Some results such as the median for wind power investment and import of green fuels are relatively close to the cost-optimal investment. Solar power and hydrogen storage are outcompeted in the cost optimum, but we see that most of the participants have opted for considerable investment into these technologies (figure 4). The participants deviate from the cost-optimal solution, and are willing to accept additional costs, if it improves other, non-monetised objectives.
@@
The resulting energy system designs of the participants show very high slack values (mean 81% and median 91%), raising questions about how well the additional cost is understood by the participants and who would bear that cost. This is striking given that almost half of the participants prioritised electricity prices when designing their system (more below). On the contrary, when the participants were explicitly asked about acceptable additional cost, the results were much lower (mean and median values of 26% and 15% respectively) in the post-study questionnaire. This stark divide between revealed and stated preferences hints at the challenge to make results more tangible, but also shows that participants expect high performance on many priorities including affordability simultaneously.
@@
4 Successful energy transitions rely on balancing trade-offs between local priorities
@@
While the aggregated results hint at the popularity of less cost-effective and less invasive technologies like solar power, heat and hydrogen storage, we can assess the participants’ choices based on their stated priorities. From the post-study questionnaire, we can see that the most commonly prioritised metrics are emissions (69%), followed by vulnerability (49.2%) and electricity price (47%). Visual impact (25.4%), system costs (24.6%), heat price (21.4%) and land use (15.9%) are not considered to be as important.
Figure 5: Obtained values from participants depending on whether a metric was prioritised; larger difference between participants who prioritised a certain metric might hint at more available options to achieve this or at higher weights of this metric. A lower value on the y-axis indicate a lower impact, which is generally desired.
@@
With the flexibility and diversity of options available, participants are forced to trade off different priorities. Most participants (57%) stated that they prioritised three or more metrics in their decision, introducing complexity into their decision-making. In fact, many participants openly acknowledged that their attempts of multi-objective planning is what they expected from policymakers too. Figure 5 evaluates whether participants using the tool achieved their stated priorities: we compare the obtained values of metrics and whether participants who stated to prioritise a metric when designing their system achieved a lower impact compared to those who did not prioritised it. Except for heat price, the group of participants who prioritised a metric, on average, identified system designs improving that metric, compared to those who did not. Some metrics, particularly costs, have larger differences than other.
@@
Figure 6: Obtained values for vulnerability and emissions from participants’ choices. Each point in a) represents a response and the according system’s vulnerability and emissions. The dashed outer line represents the complete near-optimal feasible space and the colour of a point indicates the resulting slack. The convex hulls for seven slack levels in b) show how the decision space changes with increasing slack.
@@
As participants tried to balance multiple priorities, the evaluation of how well people manage to achieve their priorities is complicated (given that their individual weighing of the different metrics is unknown to us). Figure 6a), illustrates how participants balanced the two most commonly prioritised metrics, vulnerability and emissions, and the resulting slack value through the shade of the point. Most of the participants have chosen systems with very low emissions with larger variability for vulnerability. Furthermore, figure 6b) shows how the decision space of the participants grows with increasing slack values. While it is possible to balance these two priorities, even at low slack levels, further reductions of vulnerability come at the expense of a more expensive system.
@@
Figure 7: Correlation matrix between revealed slider choice and stated priority of participants.
@@
The correlation matrix between stated priorities and revealed slider choices in figure 7 unveils patterns between the choices of the participants and their stated priorities. Participants prioritising additional costs generally reduced investment (and over-sizing of the system), whereas prioritising emissions correlates with high investment in solar power, and vulnerability with additional investment into heat storage and variable renewables. Focusing on visual impact naturally led to reduced investment in wind power (the culprit for visual impact), and instead gravitated towards elevated investments in solar power and green imports. Participants who stated a high willingness-to-pay also leaned into heat storage investment.
@@
Since the interface was restricted to feasible system configurations, participants were unable to simply minimise all of their priorities, and instead forced to make trade-offs and balance their priorities. This challenged some of the participants to think about the complexity (R01, R02, R03 in table 1), but also contributed to a feeling of being more informed. Furthermore, there were also participants who made the case for technologies and solutions that were not available in the interface, such as returning to a coal-powered energy supply (R04, R05), demand reducing measures (R06) and nuclear power (R07).